40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Nov 18 00:28:28 2010, koji, tara, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening beat.pngVCO_RCNOISE.pngcompare.pngplot_code.zipSCRN0130.zip
    Reply  Thu Nov 18 03:10:03 2010, Koji, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening IMG_3716.jpg
       Reply  Thu Nov 18 14:17:45 2010, tara, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening 
          Reply  Thu Nov 18 15:09:22 2010, Koji, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening 
             Reply  Thu Nov 18 21:42:33 2010, tara, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening Beat_VCO.png
                Reply  Thu Nov 18 22:23:20 2010, Koji, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening 
                   Reply  Fri Nov 19 01:21:53 2010, tara, DailyProgress, NoiseBudget, Noise from scattering light near the cavity's foam cap opening VCO_TF.png
Message ID: 396     Entry time: Thu Nov 18 00:28:28 2010     Reply to this: 397
Author: koji, tara 
Type: DailyProgress 
Category: NoiseBudget 
Subject: Noise from scattering light near the cavity's foam cap opening 

Today we improved some frequency noise by reducing the scattering light at the ACAV's foam cap.

 

We are searching for the cause of noise above the noise budget. 

By shaking the table and the foam box, the frequency noise seems to increase a lot at 100Hz to 500 Hz.

We are looking at both beat signal and VCO signal.

 

First, we added an ND filter just behind the EOM before the beam is split in to two paths to see if we can

reduce the power thus, reduce the scattering light. There is no significant change.

 

Next, by tapping the optics to ACAV path the noise increases significantly, so we checked and found

that the beam into the ACAV is clipped on the opening of the foam cap and causes some scattering, see [elog entry].

We tear the aluminum tape on the cap of a bit and tried to make the hole bigger, until we could not see the scattering.

The noise is lower by 1 order of magnitude around 100 Hz from beat and VCO measurements,

see fig 1 and fig 2.  We check the RCAV, which has smaller opening, but there seems to be no clipping.

We will check the scattering problem on RCAV again, by slightly remove the cap and see if there is improvement.

 

Fig 3, shows the comparison between the beat note freq and the VCO signal. A lot of mechanical peaks appears

on the beat note, but not on the VCO signal. We will want to fix the optics for beat note measurement, maybe bringing

the beam back to 3" height. Also, the beam splitter that add both beams together are very susceptible to seismic.

 

Aside from the scattering at the opening, we haven't found any major component contributing to the noise yet.

 

Unit on fig 1  (beat) is converted to Hz/ rt Hz.

Unit on fig 2 (VCO) is not converted yet, I have to look up the df/dV calibration for VCO.

 

 

 

Attachment 1: beat.png  28 kB  | Hide | Hide all
beat.png
Attachment 2: VCO_RCNOISE.png  28 kB  | Hide | Hide all
VCO_RCNOISE.png
Attachment 3: compare.png  31 kB  | Hide | Hide all
compare.png
Attachment 4: plot_code.zip  471 Bytes
Attachment 5: SCRN0130.zip  35 kB
ELOG V3.1.3-